# Do Construction Distributions Shape Formal Language Learning In German BabyLMs?

### v1

@@ 5 Construction distribution analysis
As there are no findings for the distribution of utterance-level construction in German, we conduct our own analysis using spacy. We first split larger paragraphs into individual sentences with the included senter and then annotate these with POS and dependency information. This information serves as the base of our construction annotation procedure. We devise standard construction categories in line with comparable efforts for English Cameron-Faulkner et al. (2003); Cameron-Faulkner and Noble (2013); Bunzeck and Diessel (2024), and use the spacy-annotated data to assign one of the following categories to each utterance:
@@
This holistic taxonomy is applicable to every utterance across our subcorpora. Against a manually annotated sample of 1,000 sentences, which we balanced for individual subcorpora, our classifier reaches an accuracy of approx. 95%.
Figure 1 visualizes the results of this corpus annotation process, exact proportions are reproduced in Appendix D. Generally, our results confirm earlier findings (Cameron-Faulkner et al., 2003; Cameron-Faulkner and Hickey, 2011; Cameron-Faulkner and Noble, 2013; Bunzeck and Diessel, 2024): Just like English CDS, German CDS features more questions than any other corpus, abounds with fragments, and contains comparatively few complex utterances. The Project Gutenberg data, on the other hand, is characterized by over 60% complex sentences. Interestingly, the construction distribution forms a continuum across our subcorpora. The MiniKlexikon, for example, contains considerably less complex sentences than the other written genres, but over half of its utterances are (in)transitive, canonical SV-sentences. This shows that even these particular sub-genres of child-directed linguistic input feature highly varied and specific constructional profiles that differ from each other.
@@ 6 Training data composition
We compose three different corpora of 5M words: i) one corpus maximally resembling the construction composition of child-directed speech (cds), ii) one corpus containing a drastically higher amount of complex sentences, mirroring the distribution in the Project Gutenberg data (pjg), and iii) a corpus that is averaged between these two (mix). The relative distributions of construction types can be found in Table 2.
@@
Crucially, we sample the individual utterances for our training sets from all subcorpora in our German BabyLM dataset (cf. Table 1). By doing so, we approximate a similar (if not completely equal) mixture of sources and, therefore also a similar mixture of registers, semantic content, etc. This enables us to isolate the effect of construction distributions in our model’s training data, without any interference from the possible differences between the subcorpora.
Character Subword
@@
Word-level Lexical decision 97.4% 97.6% 97.4% 84.6% 81.9% 80.8%
Surprisal 99.8% 99.8% 99.9% 91.5% 90.3% 90.1%
AntiSurprisal 99.3% 98.9% 99.7% 76.5% 75.4% 75.4%
Syntax Simple Agreement 90% 90% 95.7% 80% 84.3% 92.1%
Across a Prepositional Phrase 61.5% 65.5% 61.8% 74.8% 73.5% 75.5%
Across a Subject Relative Clause 67.1% 66% 62.4% 78.4% 73.7% 97.9%
Short Verb Phrase Coordination 69.8% 68.8% 67.9% 82.6% 93.5% 99.5%
Long Verb Phrase Coordination 53.6% 60.6% 63% 60.6% 78.8% 78%
Across Object Relative Clause 58.6% 54.2% 53% 64% 66.7% 81.6%
Within Object Relative Clause 59.8% 56.4% 72.5% 55.8% 55.7% 49.9%
Semantics XCOMPS 51.5% 49.1% 49.1% 51.4% 52% 52.3%
@@ 7 Model training and evaluation
We train small Llama models Touvron et al. (2023) with transformers Wolf et al. (2020). To account for the effect of subword tokenization, we compare character-level (3.7M parameters) and subword models (7.7M parameters) for the three datasets. We train all models for one epoch (loss curves and hyperparameters are in Appendix E) and share them on the huggingface model hub ([LINK REMOVED]). To test the effect of different random initializations and our sampling strategy, we reproduce pre-training for the cds models (see Appendix F).
In line with current best practices to linguistic probing, we use minimal pair probing datasets to evaluate our LMs’ linguistic knowledge in German. The datasets always consist of a correct/grammatical and a matched incorrect/ungrammatical string. We use minicons Misra (2022) to score the sentences. We evaluate 19 model checkpoints (10 for the first 10% of training, 9 for the remaining 90%).
@@
To gauge lexical knowledge, we adapt the experimental setup from Bunzeck et al. (2025): We use wuggy Keuleers and Brysbaert (2010) to generate 1,000 nonce words (e.g. promsen) from existing words (e.g. bremsen) and then evaluate how surprised the models are by i) the words with the context of a prepended white space (lexical decision, Le Godais et al., 2017), ii) the words in a plausible context sequence (surprisal, Hale, 2001), and iii) the words randomly inserted into implausible contexts (antisurprisal, Shafiabadi and Wisniewski, 2025). If the model is less surprised by the existing word, we count this as a correct choice in our paradigm. We calculate accuracies over the whole dataset.
@@
To evaluate our models’ semantic knowledge, we use the XCOMPS dataset He et al. (2025). It contains conceptual minimal pairs (e.g. Garnele hat einen Kopf./*Ein Bikini hat einen Kopf.) [^1] We sample 1,000 MPs with randomized replacement, as the other conditions contain implausible/wrong minimal pairs. Furthermore, the quality of translation is not optimal, as exemplified by the missing determiner in front of Garnele. that test whether LMs have acquired knowledge about conceptual properties of real-world entities. Again, we score the sentences for their likelihood and calculate accuracies for correctly rated pairs over the whole dataset.
@@ 8.1 MP probing
Table 3 shows model-wise accuracies for all minimal pair sets after training for one epoch. For the word-level evaluations, accuracy scores are generally high. Across all tasks, the character models perform with almost perfect accuracy. No effect of the constructional composition of the training data is identifiable here. For the subword models, this is not true. Here, the model trained on more questions/fragments and less complex utterances (cds) outperforms the model that approximates written language on the construction level (pjg). The improvements range from 1% for anti-surprisal to 2-3% on lexical decision.
For the syntactic tests, the picture is more nuanced. Generally speaking, all our models learn to distinguish grammatical and ungrammatical sentences with agreement phenomena. The best scores are achieved on more simplistic phenomena like simple agreement or coordination with short verb phrases. Agreement phenomena that involve longer dependencies and distracting nouns, e.g. within and across relative clauses, are the hardest to learn. For the character models, the cds model outperforms the others on three out of seven tests, including both “across subj./obj. relative clause” conditions. For three other tests, the pjg model wins out, whereas the mix model achieves the highest scores on only one test (agreement across prepositional phrases). The subword models show a somewhat different picture. Here, the pjg model wins in five categories, whereas cds and mix achieve best scores in one each.
The evaluation scores for the XCOMPS benchmark reveal that our small models do not reliably learn the conceptual knowledge underlying the included minimal pairs. Scores revolve around the chance baseline, with subword models performing slightly better than character models for 2/3 data mixtures. However, these scores are also not much worse than those reported in the XCOMPS paper (e.g. 56–69% for Llama 3.1), so more research is needed to establish ways of robust and appropriate probing of semantic knowledge in small LMs.
@@ 8.2 Learning trajectories
Figure 2 shows the learning trajectories of our models across one training epoch. In line with best practices in ML Viering and Loog (2023), we log-scale the x-axis in our plots. This allows us to also trace early learning in more detail.
@@
This pattern is also confirmed by the learning trajectories for the syntactic phenomena. While the pjg models trained on more complex utterances frequently reach the highest final scores, it is remarkable to see how the improvements for all models seem to happen in parallel. The global shape of the trajectory is the same for all syntactic tests, regardless of the construction distribution. For example, the learning curve for simple agreement is steeper for the pjg models once learning has started, but take-off points are neatly aligned. These take-off points are pushed back by the individual paradigms’ complexities – simple agreement and short VP coordination begin to improve earlier than MPs containing RCs. Finally, it is interesting to note that for the character models, word-level learning consistently stabilizes before syntactic learning, whereas both processes seem to happen concurrently in subword models (mirroring findings for English reported by Bunzeck and Zarrieß, 2025).
9 Discussion and conclusion
@@
This paper set out to investigate whether the constructional profile of CDS, which is shaped in a way to support the acquisition of functional language competence, is actually conducive to formal language learning, or whether its relative lack of complex sentences and canonical SV(X) utterances makes it so impoverished that meaningful formal learning does not happen. The results of our utterance-level corpus analysis for German align with earlier findings on CDS and book language for English Cameron-Faulkner et al. (2003); Cameron-Faulkner and Noble (2013); Bunzeck and Diessel (2024) and Irish Cameron-Faulkner and Hickey (2011), adding to the growing evidence that this linguistic distribution is fairly universal, at least in WEIRD societies Henrich (2024).
@@
The results of our language modeling experiments are much more surprising. The constructional profile of training data is not overly important for training small, developmentally plausible LMs from scratch. Rather, starting/turning points of the resulting learning trajectories are mostly determined by the respective amount of training steps. Despite models trained with more complex input resulting in slightly better performance, they do not begin to learn earlier. Global learning trajectories are extremely similar, only the local magnitude differs between different constructional setups. For word-level learning processes such as lexical decision or (anti)surprisal tests, data with more fragments and questions even seems to be rather conducive, whereas models trained on more complex utterances do not learn this level of knowledge equally well.
@@
What does this now mean for theories of language acquisition? On a formal level, there seems to be no disadvantage for models trained on allegedly less “complex” or somewhat impoverished data. Despite more complex data leading to slightly better benchmark scores, the learning trajectories remain largely unaffected. What really shapes the learning process is the amount of input, not its formal complexity (similar to findings for children by Huttenlocher et al., 1991; Rowe, 2012). An increase in appropriate construction types for child-rearing and functional language learning (like questions, imperatives, or fragments) does not hinder formal learning (if only reduce its magnitude slightly). Conversely, it even enables word-level learning to converge to a better end state. This also aligns with a broader trend found in language acquisition studies — the complexity and quality of input can indeed predict later language skills Noble et al. (2020); Alroqi et al. (2023), but the ground level is always extremely high already: being a competent user of the language itself. Furthermore, quality varies with many more extralinguistic factors like the number of siblings Laing and Bergelson (2024) or cultural factors Bergelson et al. (2023); Bunce et al. (2024).
@@
Finally, our findings also add to the growing body of research on BabyLMs Warstadt et al. (2023); Hu et al. (2024). Similarly to English models, our German BabyLMs only need little data – the cds dataset contains approx. 820,000 sentences, and given the estimation by Cameron-Faulkner et al. (2003) that children hear around 7,000 utterances per day, our data approximates the number of utterances heard over only 120 days – to learn a fair amount of syntax and almost impeccable lexical knowledge, with trajectories mirroring those of English models Bunzeck and Zarrieß (2025). We hope that our dataset enables other scholars to carry out experiments with developmentally plausible LMs beyond the dominating English LMs, and that our data sources provide inspiration to those compiling BabyLM corpora for their own languages of interest.
Appendix A Limitations and ethical considerations
A.1 Limitations
@@
Our study is limited by data availability. Creating a full-fledged 100M-token BabyLM dataset is currently out of question, as neither CHILDES nor other sources contain even remotely enough data for languages other than English. Principally, synthetic corpora like the TinyStories dataset Eldan and Li (2023), which contains children’s stories generated by GPT-3 or TinyDialogues by Feng et al. (2024) would provide an unlimited source of training data. However, our inspection of their generated dialogues yielded that they drastically underestimate the high numbers of grammatical fragments, questions and short SV(X)-utterances in real-world data. Similarly, there are little to no evaluation sets for German beyond those that we included/creates ourselves, especially on the syntactic level.
@@
Moreover, actual developmental plausibility also hinges on the inclusion of other modalities. For audio data, there are few CHILDES subcorpora and other corpora that contain phonetic information Lavechin et al. (2023), but larger models need to be trained on more data, e.g. audiobooks Lavechin et al. (2025). A middle ground is training on textual phonetic transcriptions generated from raw text, e.g. for the BabyLM data Goriely et al. (2024); Bunzeck et al. (2025). More recently, also video recordings from infant-mounted cameras have been used to train on combined visual and auditory input modalities Wang et al. (2023); Vong et al. (2024); Long et al. (2024). The inclusion of such data could help to disentangle learning processes further.
A.2 Ethical considerations
@@
Given the nature of this work, there are no specific ethical concerns to address. However, we want to emphasize that BabyLMs are not actual babies, but rather abstractions, or models in the original scientific sense, of the distributional, frequency-driven aspects of their learning capacity. All claims regarding their implications for language development in the real world should be understood in this context, which we also attempted to explicate by distinguishing functional and formal aspects of learning.
Appendix B Excluded corpora

### v2

@@ 5 Construction distribution analysis
As there are no findings on the distribution of utterance-level construction in German, we conduct our own analysis using spacy Honnibal et al. (2020). We first split larger paragraphs into individual sentences with the included senter and then annotate these with POS and dependency information. This information serves as the base of our construction annotation procedure. We devise standard construction categories in line with comparable efforts for English Cameron-Faulkner et al. (2003); Cameron-Faulkner and Noble (2013); Bunzeck and Diessel (2024), and assign one of the following categories to each utterance:
@@
This holistic taxonomy is applicable to every utterance in our corpus. For a balanced, manually annotated sample of 1,000 sentences our classifier reaches an accuracy of approx. 95%.
Figure 2 visualizes the results of this annotation process, exact proportions are reproduced in Appendix C. Generally, our results confirm earlier findings (Cameron-Faulkner et al., 2003; Cameron-Faulkner and Hickey, 2011; Cameron-Faulkner and Noble, 2013; Bunzeck and Diessel, 2024): Just like English CDS, German CDS features more questions than any other corpus, abounds with fragments, and contains comparatively few complex utterances. The Project Gutenberg data, on the other hand, is characterized by over 60% complex sentences. Interestingly, the construction distribution forms a continuum across our subcorpora. The MiniKlexikon, for example, contains considerably less complex sentences than the other written genres, but over half of its utterances are (in)transitive, canonical SV-sentences. This shows that even these particular sub-genres of child-directed linguistic input feature highly varied and specific constructional profiles that differ from each other.
@@ 6 Training data composition
We compose three different corpora of 5M words: (1) one corpus maximally resembling the construction composition of child-directed speech (cds), (2) one corpus containing a drastically higher amount of complex sentences, mirroring the distribution in the Project Gutenberg data (pjg), and (3) a corpus that is averaged between these two (mix). The relative distributions of construction types can be found in Table 2.
@@
Crucially, we sample the individual utterances for our training sets from all subcorpora in our German BabyLM dataset. By doing so, we approximate a similar (if not completely equal) mixture of sources and, therefore also a similar mixture of registers, semantic content, etc. This enables us to isolate the effect of construction distributions in our model’s training data, without any interference from the possible differences between the subcorpora.
Character Subword Llama 3.2 1B
@@
Word-level Lexical decision 97.4% 97.6% 97.4% 84.6% 81.9% 80.8% 69.6%
Surprisal 99.8% 99.8% 99.9% 91.5% 90.3% 90.1% 98%
AntiSurprisal 99.3% 98.9% 99.7% 76.5% 75.4% 75.4% 87.4%
Syntax Simple Agreement 90% 90% 95.7% 80% 84.3% 92.1% 95.71%
Across a Prepositional Phrase 61.5% 65.5% 61.8% 74.8% 73.5% 75.5% 83%
Across a Subject Relative Clause 67.1% 66% 62.4% 78.4% 73.7% 97.9% 99.7%
Short Verb Phrase Coordination 69.8% 68.8% 67.9% 82.6% 93.5% 99.5% 99.9%
Long Verb Phrase Coordination 53.6% 60.6% 63% 60.6% 78.8% 78% 90.5%
Across Object Relative Clause 58.6% 54.2% 53% 64% 66.7% 81.6% 86.1%
Within Object Relative Clause 59.8% 56.4% 72.5% 55.8% 55.7% 49.9% 61.4%
Semantics XCOMPS 51.5% 49.1% 49.1% 51.4% 52% 52.3% 58.9%
@@ 7 Model training and evaluation
We train small Llama models Touvron et al. (2023) with transformers Wolf et al. (2020). To account for the effect of subword tokenization, we compare character-level (3.7M parameters) and subword models (7.7M parameters) for the three datasets. We train all models for one epoch (loss curves and hyperparameters are in Appendix D) and share them on Hugging Face. [^2] https://huggingface.co/collections/bbunzeck/german-babylm-67b868e08ff8782a9814ceaf To test the effect of different random initializations and our sampling strategy, we reproduce pre-training for the cds models (see Appendix E).
In line with current best practices to linguistic probing, we use minimal pair datasets to evaluate our LMs’ linguistic knowledge in German. The datasets always consist of a correct/grammatical and a matched incorrect/ungrammatical string. We use minicons Misra (2022) to score the sentences and evaluate 19 model checkpoints per model (10 for the first 10% of training, 9 for the remaining 90%). As an additional ablation, we also evaluate the multilingual Llama 3.2 1B [^3] https://huggingface.co/meta-llama/Llama-3.2-1B on all probing paradigms. Currently, no monolingual German Llama models exist. Therefore, the medium-sized 1B-parameter version of Llama 3.2, which is trained on a considerable amount of German language data, is a useful baseline for expected benchmark scores enabled through a higher model capacity and more training data.
@@
Language acquisition first involves learning what words are, i.e. which (sound) sequences map to word-level items in the mental lexicon, before learning how they combine. To gauge this most basic learning step, we adapt the experimental setup from Bunzeck et al. (2025): We use wuggy Keuleers and Brysbaert (2010) to generate 1,000 nonce words (e.g. promsen) from existing words (e.g. bremsen) and then evaluate how surprised the models are by (1) the words with the context of a prepended white space (lexical decision, Le Godais et al., 2017), (2) the words in a plausible context sequence (surprisal, Hale, 2001), and (3) the words randomly inserted into implausible contexts (antisurprisal, Shafiabadi and Wisniewski, 2025). If the model is less surprised by the existing word, we count this as a correct choice in our paradigm. We calculate accuracies over the whole dataset.
@@
To evaluate our models’ semantic knowledge, we use the XCOMPS dataset He et al. (2025). It contains conceptual minimal pairs (e.g. Garnele hat einen Kopf./*Ein Bikini hat einen Kopf.) [^4] We sample 1,000 MPs with randomized replacement, as the other conditions contain implausible/wrong minimal pairs. Furthermore, the quality of translation is not optimal, as exemplified by the missing determiner in front of Garnele. that test whether LMs have acquired knowledge about conceptual properties of real-world entities. Again, we score the sentences for likelihood and calculate accuracy over the whole dataset.
@@ 8.1 MP probing
Table 3 shows model-wise accuracies for all minimal pair sets after training for one epoch. For the word-level evaluations, accuracy scores are generally high. Across all tasks, the character models perform with almost perfect accuracy. No effect of the constructional composition of the training data is identifiable here. For the subword models, this is not true. Here, the model trained on more questions/fragments and less complex utterances (cds) outperforms the model that approximates written language on the construction level (pjg). The improvements range from 1% for anti-surprisal to 2-3% on lexical decision. Interestingly, the very large ablation model (Llama 3.2 1B) performs the worst on isolated lexical decision, but reaches high scores in the surprisal setting.
For the syntactic tests, the picture is more nuanced. Generally speaking, all our models learn to distinguish most types of grammatical and ungrammatical sentences involving agreement phenomena. The best scores are achieved on more simplistic phenomena like simple agreement or coordination with short verb phrases. Agreement phenomena that involve longer dependencies and distracting nouns, e.g. within and across relative clauses, are the hardest to learn. For the character models, the cds model outperforms the others on three out of seven tests, including both “across subj./obj. relative clause” conditions. For three other tests, the pjg model wins out, whereas the mix model achieves the highest scores on only one test (agreement across prepositional phrases). It should be noted, that for most phenomena, the character models do perform well above chance (by a margin of 10–20%), but still frequently make errors. The subword models show a somewhat different picture, with scores being generally higher and approximating perfect performance on 3/7 phenomena. Regarding construction distributions, the pjg model wins in five categories, whereas cds and mix only achieve best scores in one each. Here, the 1B-parameter Llama model outperforms our BabyLMs on 5/7 phenomena.
The scores on XCOMPS reveal that our small models do not reliably learn the conceptual knowledge underlying the included minimal pairs. Scores revolve around the chance baseline, with subword models performing slightly better than character models for 2/3 data mixtures. Nonetheless, these scores are also not considerably worse than the performance of our ablation model (58.9%).
@@ 8.2 Learning trajectories
Figure 3 shows the learning trajectories of our models across one training epoch. As there are no intermediate checkpoints available for the 1B-parameter ablation model, we only report trajectories for our self-trained models. In line with best practices in ML Viering and Loog (2023), we log-scale the x-axis in our plots. This allows us to also trace early learning in more detail.
@@
This pattern is also confirmed by the learning trajectories for the syntactic phenomena. While the pjg models trained on more complex utterances frequently reach the highest final scores, it is remarkable to see how the improvements for all models seem to happen in parallel. The global shape of the trajectory is the same for all syntactic tests, regardless of the construction distribution. For example, the learning curve for simple agreement is steeper for the pjg models once learning has started, but take-off points are neatly aligned. These take-off points are pushed back by the individual paradigms’ complexities — simple agreement and short VP coordination begin to improve earlier than MPs containing RCs. Finally, it is interesting to note that for the character models, word-level learning consistently stabilizes before syntactic learning, whereas both processes seem to happen concurrently in subword models (mirroring findings for English, cf. Bunzeck and Zarrieß, 2025). As our models do not learn to distinguish the semantic minimal pairs, the corresponding learning curves remain flat and performance differences are likely due to chance.
9 Discussion
@@
This paper set out to investigate whether the constructional profile of CDS, which is shaped in a way to support the acquisition of functional language competence, actually influences LMs’ formal language learning, and whether its relative lack of complex sentences and canonical SV(X) utterances makes it less useful training data, or too “impoverished” for meaningful formal learning to happen. The results of our utterance-level corpus analysis for German align with earlier findings on CDS and book language for English Cameron-Faulkner et al. (2003); Cameron-Faulkner and Noble (2013); Bunzeck and Diessel (2024) and Irish Cameron-Faulkner and Hickey (2011), adding to the growing evidence that this linguistic distribution is fairly universal, at least in WEIRD societies Henrich (2024).
@@
From a language modeling perspective, the constructional profile of training data is not overly important for the resulting performance on linguistic benchmarks. Rather, starting/turning points of the resulting learning trajectories are mostly determined by the respective amount of training steps. Despite models trained with more complex input resulting in slightly better performance, they do not begin to learn earlier. Global learning trajectories are extremely similar, only the local magnitude differs between different constructional setups. This provides further evidence that LMs based on the Transformer architecture Vaswani et al. (2017) not only memorize language from their training data, but generalize to the underlying patterns. The same holds true or word-level learning processes such as lexical decision or (anti)surprisal tests, where data with more fragments and questions even seems to be rather beneficial. Furthermore, the comparison of our results to the Llama 3.2 1B model shows that rather high scores are already attainable with small models and little data (only on long VP-coordination do our models underperform).
@@
What does this now mean for theories of language acquisition? This study was inspired by findings of construction-based corpus analyses Cameron-Faulkner et al. (2003); Cameron-Faulkner and Hickey (2011); Bunzeck and Diessel (2024), which argue that the specific constructional profile of CDS is beneficial to acquisition. Of course, LMs and minimal pair evaluations do not directly correspond to the learning processes in humans and we cannot make causal claims about them. Yet, our methodology can provide evidence as to what kinds of input data is beneficial to a purely statistical learner (that does not even tap into the functional side of language, cf. Mahowald et al., 2024), an abstraction that is highly relevant to usage-based theories Ambridge et al. (2015). On a formal level, there seem to be comparatively little disadvantages for models trained on less “complex” or somewhat impoverished data. Despite more complex data leading to slightly better benchmark scores, the learning trajectories remain largely unaffected (although somewhat erratic, cf. Bunzeck and Zarrieß, 2024). What really shapes the learning process in our LMs is the amount of input, not its formal complexity (similar to findings for children by Huttenlocher et al., 1991; Rowe, 2012). An increase in appropriate construction types for child-rearing (like questions, imperatives, or fragments) does not hinder formal learning (if only reduce its magnitude slightly). As CLAMS only focuses on subject-verb agreement in canonical SV(X)-sentences, it is rather surprising that the much higher amount of questions in the cds dataset does not negatively affect performance, although the subjects’ and predicates’ positions are switched in German yes/no-questions. Conversely, the cds dataset even enables word-level learning to converge to a better end state. This also aligns with a broader trend found in language acquisition studies — the complexity and quality of input can indeed predict later language skills Noble et al. (2020); Alroqi et al. (2023), but the ground level is always extremely high already: being a competent user of the language itself. Furthermore, quality varies with many more extralinguistic factors like the number of siblings Laing and Bergelson (2024) or cultural factors Bergelson et al. (2023); Bunce et al. (2024).
10 Conclusion
@@
Our findings add to the growing body of research on BabyLMs Warstadt et al. (2023); Hu et al. (2024). Similarly to English models, our German BabyLMs only need little data — the cds dataset contains approx. 820,000 sentences, and given the estimation by Cameron-Faulkner et al. (2003) that children hear around 7,000 utterances per day, our data approximates the number of utterances heard over only 120 days — to learn a fair amount of syntax and almost impeccable lexical knowledge, with trajectories mirroring those of English models Bunzeck and Zarrieß (2025). We hope that our dataset enables other scholars to carry out experiments with developmentally plausible LMs beyond the dominating English LMs, and that our data provides inspiration to those compiling BabyLM corpora for other languages.
Limitations
@@
Our study is limited by data availability. Creating a full-fledged 100M-token BabyLM dataset with only child-directed speech or other explicitly child-directed materials is currently out of question, as neither CHILDES nor other sources contain even remotely enough data for languages other than English. To reach higher token counts, padding with larger data sets, e.g. more tokens from the OpenSubtitles dataset, would be necessary. Principally, synthetic corpora like the TinyStories dataset Eldan and Li (2023), which contains children’s stories generated by GPT-3 or TinyDialogues by Feng et al. (2024) would provide an unlimited source of training data. However, our inspection of their generated dialogues yielded that they drastically underestimate the high numbers of grammatical fragments, questions and short SV(X)-utterances in real-world data. Similarly, there are little to no evaluation sets specifically aimed at German, beyond those that we included/creates ourselves, especially on the syntactic level. Only very recently, evaluation datasets like the massively multilingual MultiBLiMP have begun to fill this gap Jumelet et al. (2025). Also, such minimal pair datasets are principally at odds with the usage-based, constructionist view on language development, because they are grounded in the Generativist notion of defining rules that can determine whether an utterance belongs to a language or not, whereas usage-based linguistics has adopted a network-based, associative model of linguistic knowledge Diessel (2019, 2023). As of late, these developments have begun to make their way into the broader LM evaluation landscape Weissweiler et al. (2025), and novel evaluation methods like measuring affinities between lexical items and testing if different constructions manifest from them Rozner et al. (2025a, b) provide promising future research avenues.
@@
Moreover, actual developmental plausibility also hinges on the inclusion of other modalities. For audio data, there are few CHILDES subcorpora and other corpora that contain phonetic information Lavechin et al. (2023), but larger models need to be trained on more data, e.g. audiobooks Lavechin et al. (2025). A middle ground is training on textual phonetic transcriptions generated from raw text, e.g. for the BabyLM data Goriely et al. (2024); Bunzeck et al. (2025); Goriely and Buttery (2025b). More recently, also video recordings from infant-mounted cameras have been used to train on combined visual and auditory input modalities Wang et al. (2023); Vong et al. (2024); Long et al. (2024). The inclusion of such data could help to disentangle learning processes further.
Ethical considerations
@@
Given the nature of this work, there are no specific ethical concerns to address. However, we would like to stress that, of course, BabyLMs are not supposed to simulate real babies, but rather to instantiate abstractions, or models in the original scientific sense, of the distributional, frequency-driven aspects of their learning capacity. All claims regarding their implications for language development in the real world should be understood in this context, which we also attempted to explicate by distinguishing functional and formal aspects of learning.
Acknowledgments
This research has been funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) — CRC-1646, project number 512393437, project A02.
Appendix A Excluded corpora